An AC circuit includes a capacitor of 2.0 micro ferads and an inductor of inductance 8.3 milihenerry. The AC voltage supply is set at a frequency of f= 100hz.

A) what is the capacitive reactance? B) what is the inductive resistance?
Answer #1

Given data,

C = 2 F = 2 * 10-6 F

L = 8.3 mH = 8.3 * 10-3 H

f = 100 hz

Formula required to solve the problem,

XL = L = 2 f L               [Where XL= Inductive Reactance in Ohm, f = Frequency in hz, L = Inductance in Henry]

XC = 1/(C) = 1/ (2 f C)   [Where XC= Capacitive Reactance in Ohm, f = Frequency in hz, C = Capacitance in Farad]

  • Capacitive Reactance :-

XC = 1/(C) = 1/ (2 f C) = 1/ (2 *100* 2 * 10-6) = 795.774

  • Inductive Reactance :-

XL = L = 2 f L = 2 * 100 * 8.3 * 10-3 = 5.215